OpenStudy (anonymous):

Aidan drives to school and back each day. The school is 16 miles from his home. He averages 40 miles per hour on his way to school. If his total trip takes 1 hour, at approximately what speed does Aidan drive home?

OpenStudy (mathmale):

speed = distance/time , and, conversely, time = distance/ speed. How long does it take Aidan to drive to school? With this info, how do you answer the second part of the question, namely, determine the speed at which Aidan drives home?

OpenStudy (anonymous):

1=16/40 ?

OpenStudy (anonymous):

would that solve how long it takes him to go to school?

jigglypuff314 (jigglypuff314):

the amount of time it take for him to go from home to school is 16/40 = then the amount of time it took for him to get back home is 1 - (16/40) = but want the average speed so...

jigglypuff314 (jigglypuff314):

average speed going back home = (distance) / (amount of time)

OpenStudy (anonymous):

so just 16/.06?

jigglypuff314 (jigglypuff314):

if you meant 16/0.6 then yeah and it should give you a number :)

OpenStudy (mathmale):

We assume that Aidan travels at one speed going to school and at a different speed coming home. The total amount of time spent en route is 1 hour. If use the formula distance = rate * time, and solve for time, we get time =distance/rate. Here the distance is 16 miles and the rate at which A. drives to school is 40 mph. How long does it take Aidan to drive to school? Next: subtract that time from 1 hour. What's left? That's the amount of time Aidan spends driving home. How do we figure out Aidan's average speed on the way home?

OpenStudy (mathmale):

You're on the right track, and I should have realized that earlier. Take jigglypuff314's suggestion: (rate at which Aidan drives home) = (16 miles) / (0.6 hour). Can you simplify this to obtain the number of miles per hour (which is the goal of this problem)?